x(t) hết) WWW.HBHB (in) (b)

Consider two periodic signals, please

(a) determine and draw the output of y(t)=x(2t)*h(t) if T >8.

(b) Try to figure out and draw the frequency responses of x(t), h(t), and y(t).
Answer #1

* Given the 819nal, X40) -25 -1/41114 alt) - Time period T. 1991 x(2+2 Time period (772). alt+14) 21 (24-T): 1. Т X (27) A -T711) = xet) * h 4 ) 494) = {-1814<*: < 1/4 +774 of T2 21 L-112. 715 -37/2 -1/2 Th 37/2 6 Abone the solution we have draw the